Treating Drinking Water Phosphate can be removed from drinking-water supplies by treating the water with Ca(OH)2. How much Ca(OH), is required to remove 90% of the PO43-from 4.5 × 106 L of drinking water containing 25 mg/L of PO
Answer #1

Given

Concentration of PO4 3- = 25 mg/L = 25*10^-3 g / L

1 L of drinking water contains = 25 x 10^-3 g of PO43-

4.5 X 10^6 L of drinking water contains = 25 x 10^-3 x 4.5 x10^6 = 112500 g

remove 90 % of PO43-

PO43- removed = 0.90 x 112500 = 101250 g

moles of PO43- removed = mass / molecular weight

= 101250/95

= 1065.789 moles

The balanced chemical equation

5 Ca(OH2) + 3 PO4 3- = Ca5OH(PO4)3 +9 OH-

From the stoichiometry of the reaction

3 mole of PO43- = 5 moles of Ca(OH)2

1065.789 moles of PO43- = (5/3) x 1065.789 = 1776.316 moles of Ca(OH)2

molecular mass of Ca(OH)2 = 74.1 g /mol

Mass of Ca(OH)2 recquired = moles x molecular weight

= 1776.316 moles x 74 g/mol

= 131447.368 g